You have been asked to prepare a December cash budget for Ashton Company, a distributor of exercise equipment. The following information is available about the company’s operations:

  1. The cash balance on December 1 is $41,000.

  2. Actual sales for October and November and expected sales for December are as follows:

October November December
Cash sales $ 68,400 $ 87,400 $ 95,800
Sales on account $ 415,000 $ 614,000 $ 692,000

Sales on account are collected over a three-month period as follows: 20% collected in the month of sale, 60% collected in the month following sale, and 18% collected in the second month following sale. The remaining 2% is uncollectible.

  1. Purchases of inventory will total $376,000 for December. Thirty percent of a month’s inventory purchases are paid during the month of purchase. The accounts payable remaining from November’s inventory purchases total $174,500, all of which will be paid in December.

  2. Selling and administrative expenses are budgeted at $452,000 for December. Of this amount, $63,500 is for depreciation.

  3. A new web server for the Marketing Department costing $86,500 will be purchased for cash during December, and dividends totaling $10,500 will be paid during the month.

  4. The company maintains a minimum cash balance of $20,000. An open line of credit is available from the company’s bank to increase its cash balance as needed.

Required:

1. Calculate the expected cash collections for December.

2. Calculate the expected cash disbursements for merchandise purchases for December.

3. Prepare a cash budget for December. Indicate in the financing section any borrowing that will be needed during the month. Assume that any interest will not be paid until the following month

Complete this question by entering your answers in the tabs below.

  • Req 1

1. Calculate the expected cash collections for December.

2. Calculate the expected cash disbursements for merchandise purchases for December.

Prepare a cash budget for December. Indicate in the financing section any borrowing that will be needed during the month. Assume that any interest will not be paid until the following month.

Ashton Company
Cash Budget
For the Month of December
Beginning cash balance
Add collections from customers
Total cash available
Less cash disbursements:
Payments to suppliers for inventory
Selling and administrative expenses
New web server
Dividends paid
Total cash disbursements
Excess (deficiency) of cash available over disbursements
Financing:
Borrowings
Repayments
Interest
Total financing
Answer #1

1.) Expected Cash Collections for December December Cash Sales + December sales on account x 20% + November Sales on account

$ 41,000 $6,77,300 $7,18,300 Ashton Company Cash Budget For The Month of December Beginning Cash Balance Add: Collections fro

1.) Expected Cash Collections for December = December Cash Sales + December sales on account x 20% + November Sales on account x 60% + October Sales on account x 18% = 95,800 + 692,000 x 20% +614,000 x 60% +415,000 x 18% = $ 6,77,300 2.) Expected Cash Disbursements for Merchandise purchases for December - December Purchases x 30% + November Accounts Payable = 376,000 x 30% + 174,500 = $ 2,87,300

$ 41,000 $6,77,300 $7,18,300 Ashton Company Cash Budget For The Month of December Beginning Cash Balance Add: Collections from Customers Total Cash Available Less: Cash Disbursements: Payments to Suppliers for Inventory Selling and Administrative Expenses New Web Server Dividends Paid Total Cash Disbursements Excess(Deficiency of Cash available over disbursements Financing: Borrowings Repayments Interest Total Financing Ending Cash Balance $2,87,300 $3,88,500 $ 86,500 $ 10,500 $ 7,72,800 $ -54,500 $ 74,500 $ 20,000 Selling and Administrative Expenses paid = Selling and Administrative Expenses - Depreciation = 452,000 - 63,500 = $ 388,500 Borrowings = Deficiency of Cash + Desired Ending Cash Balance = 54,500 + 20,000 = $ 74,500